# Unlocking Doping Effects on Altermagnetism in MnTe: Emergence of Quasi-altermagnetism

### v2

@@ IV Summary
For paired non-magnetic substitutions, the supercell produces 120 possible configurations, and based on their space group symmetries, they are segregated into five different families. We discovered that perfect altermagnetism persists in three families covering $\approx$ 46% of all possible configurations, while in the remaining two, the opposite-spin sublattices are not rotational-symmetry related and thus do not qualify as ideal altermagnets. However, their band structures retain the key altermagnetic features, including momentum-dependent spin splitting and nearly compensated net magnetization. For isovalent doping, deviations from perfect altermagnetism are minimal, whereas hole and electron doping produce more pronounced deviations, but still result in near-zero magnetization due to alternating spin dominance across the Brillouin zone. We have termed this behaviour as quasi-altermagnetism, characterized by significant momentum-dependent spin splitting with a small residual moment. Moreover, through model Hamiltonians, we investigated the microscopic origin of the quasi-altermagnetism and identified the perturbations that lead to the transition from ideal altermagnetism to quasi-altermagnetism.
@@
In summary, our results identify the general conditions under which altermagnetism persists or deviates under non-magnetic doping, providing a realistic assessment of robustness altermagnets for spintronic applications. Even when ideal symmetry conditions are broken, MnTe retains essential altermagnetic characteristics, and suitable dopant selection offers a viable pathway for tailoring its properties for device integration. We hope that our predictions could be experimentally tested in the near future.
@@ Acknowledgments
Author contributions
N.D., A.N., and B.R.K.N. conceived the problem statements and designed the project. N.D. performed the DFT calculations. Symmetry analysis was carried out by A.B., N.D., A.D., A.M., and M.A.R. AHE study was performed by N.D., A.B., and M.A.R. The model Hamiltonian study was done by A.M. and B. R. K. N. All authors contributed to the analysis of the results and the preparation of the manuscript. The project was overall supervised by A. N. and B. R. K. N.
Competing interests
The authors declare no competing interests.
@@ Appendix A Altermagnetism in primitive cell of MnTe with paired non-magnetic atom substitutions
Figure 10: Hexagonal supercell and orthorhombic primitive cell of pair non-magnetic atom doped MnTe with MSG: $Pm^{\prime}m^{\prime}2$ (No.25.60). Panel (a) shows the hexagonal MnTe supercell with a pair of non-magnetic atom substitutions, exhibiting MSG $Pm^{\prime}m^{\prime}2$ (No.25.60). The corresponding constant energy surfaces for the spin-up and spin-down channels at E=EF-0.7 eV are shown in panel (c), revealing $g$-wave altermagnetic characteristics. Panel (b) depicts the primitive cell corresponding to the hexagonal structure in (a), and its constant energy surface at E=EF-0.7 eV is presented in panel (d). This structure exhibits $d$-wave altermagnetism. Panel (e) and (f) correspond to the Brillouin zones of hexagonal and orthorhombic structures and (g) illustrates how three orthorhombic Brillouin zones, colored blue, green, and orange, combine to form a hexagonal Brillouin zone outlined in black. It can be observed that the nodal planes of the three orthorhombic structures collectively give rise to four nodal planes in the hexagonal Brillouin zone, leading to the emergence of $g$-wave altermagnetism.
Fig. 10 (a) shows the hexagonal MnTe supercell with a pair of non-magnetic atom substitutions, exhibiting MSG $Pm^{\prime}m^{\prime}2$ (No. 25.60). The constant energy surfaces for the spin-up and spin-down channels, computed at E = EF – 0.7 eV, are shown in Fig. 10(c). These surfaces reveal a $g$-wave altermagnetic character, characterized by distinct yet symmetric features connected through a sixfold roto-invertion symmetry in each spin channel.
@@
Appendix B Symmetry operations connecting opposite spin sublattices in doped MnTe structures
For the single dopant scenario, the system consistently exhibits altermagnetic ordering. In contrast, the paired dopant scenario leads to several distinct configurations [see Fig. 3(e) of main text], with 46.67% maintaining altermagnetism, while the rest do not. Table 1 summarizes the symmetry operations connecting the opposite spin sublattices and corresponding magnetic classes for all these doped systems. For MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211), $Fm^{\prime}m^{\prime}2$ (No.42.222), and $Pm^{\prime}m^{\prime}2$ (No.25.60) the opposite spin sub-lattices are connected by symmetries other than translation or inversion, therefore being classified as altermagnets. On the other hand, for $P\bar{3}m^{\prime}1$ (No.164.89) and $C2^{\prime}/m^{\prime}2$ (No.12.62), which we refer to as quasi-altermagnets, there exist no symmetry operations that connect the up and down spin sublattices. This absence results in a lack of compensation between the spin channels, thereby lifting the degeneracy between the up and down spin bands throughout the Brillouin zone.
Table 1: Symmetry operations connecting opposite spin sublattices and AHC components in doped MnTe. Magnetic space groups (MSGs) obtained from single and pair doping in MnTe, along with the symmetry operations connecting opposite spin sublattices, their corresponding magnetic classes (AM: altermagnet, Q-AM: quasi-altermagnet), and the resulting anomalous Hall conductivity (AHC) components.
MSG Symmetry between up and down spins Magnetic class AHC component
187.211 $S^{\pm}_{6z}$, $M_{z}$ AM None
164.89 No symmetry Q-AM $\sigma_{xy}$
42.222 $C_{2z}$, $M_{y}$ AM $\sigma_{xy}$
25.60 $C_{2z}$, $M_{x}$ AM $\sigma_{xy}$
12.62 No symmetry Q-AM $\sigma_{xy}$, $\sigma_{yz}$
Appendix C Effect of magnetic atom doping in MnTe
@@
A single substitution, as expected with different local spin moments at Fe site, breaks the rotational symmetry connecting the opposite spin sublattices, resulting in MSG $P3m^{\prime}1$ (No.156.51). The corresponding crystal structure and band structure are shown in Fig. 11 (a) and (d). The bands exhibits the characteristics of quasi-altermagnetism. In addition, new defect states are introduced, as shown in the band structure (Fig. 11 (d)).
@@
We further examined the pair substitution cases, where two Mn atoms with opposite spin moments are substituted by Fe. Similar to the pair of non-magnetic atom substitutions discussed in the main text, a pair of magnetic atom substitutions also leads to multiple configurations with different MSG. When the dopant positions are arranged in such a way that mirror symmetry (Mz) of the system is preserved, then the resulting structure belongs to MSG $P\bar{6}^{\prime}m^{\prime}2$ (No.187.211) and will exhibit altermagnetic characteristics. The other configurations, where the proper or improper rotational symmetry is absent, lead to quasi-altermagnetism. Compared to non-magnetic atom substitution the probability of retaining altermagnetism is less when multiple magnetic atoms are substituted, since the magnetic neutrality is also affected. The other cases will result into quasi-altermagnetic conditions. Band structures displaying quasi-altermagnetic characteristics and ideal altermagnetic characteristics are shown in Fig. 11 (e) and (f). These results demonstrate that the quasi-altermagnetism discussed in the main text for non-magnetic atom doping is not restricted to such cases; it can also occur when magnetic atoms are introduced.

### v3

@@ IV Summary
For paired non-magnetic substitutions, the supercell produces 120 possible configurations, and based on their SG symmetries, they are segregated into five different families. We discovered that perfect altermagnetism persists in three families covering $\approx$ 46% of all possible configurations, while in the remaining two, the opposite-spin sublattices are not rotational-symmetry related and thus do not qualify as ideal altermagnets. However, their band structures retain the key altermagnetic features, including momentum-dependent spin splitting and nearly compensated net magnetization. For isovalent doping, deviations from perfect altermagnetism are minimal, whereas hole and electron doping produce more pronounced deviations, but still result in near-zero magnetization due to alternating spin dominance across the Brillouin zone. We have termed this behavior as quasi-altermagnetism, characterized by significant momentum-dependent spin splitting with a small residual moment. Moreover, through model Hamiltonians, we investigated the microscopic origin of the quasi-altermagnetism and identified the perturbations that lead to the transition from ideal altermagnetism to quasi-altermagnetism.
@@
In summary, our results identify the general conditions under which altermagnetism persists or deviates under non-magnetic doping, providing a realistic assessment of robustness altermagnets for spintronic applications. Even when ideal symmetry conditions are broken, MnTe retains essential altermagnetic characteristics, and suitable dopant selection offers a viable pathway for tailoring its properties for device integration. The quasi-altermagnetic phase proposed in this work, characterized by nonidentical spin splitting and Fermi surface topology, may enable other intriguing spin-dependent transport properties, such as the spin Hall effect and spin transfer torques. These will be interesting avenues for future work. We hope that our predictions could be experimentally tested in the near future.
@@ Appendix A Altermagnetism in primitive cell of MnTe with paired non-magnetic atom substitutions
Figure 10: Hexagonal supercell and orthorhombic primitive cell of pair non-magnetic atom doped MnTe with SG $Pmm2$ (No. 25). Panel (a) shows the hexagonal MnTe supercell with a pair of non-magnetic atom substitutions, exhibiting SG $Pmm2$ (No. 25). The corresponding constant energy surfaces for the spin-up and spin-down channels at E=EF-0.7 eV are shown in panel (c), revealing $g$-wave altermagnetic characteristics. Panel (b) depicts the primitive cell corresponding to the hexagonal structure in (a), and its constant energy surface at E=EF-0.7 eV is presented in panel (d). This structure exhibits $d$-wave altermagnetism. Panel (e) and (f) correspond to the Brillouin zones of hexagonal and orthorhombic structures and (g) illustrates how three orthorhombic Brillouin zones, colored blue, green, and orange, combine to form a hexagonal Brillouin zone outlined in black. It can be observed that the nodal planes of the three orthorhombic structures collectively give rise to four nodal planes in the hexagonal Brillouin zone, leading to the emergence of $g$-wave altermagnetism.
Figure 12: Schematic illustration of the band diagram representing different classes of collinear magnetic systems. For antiferromagnets/ ferrimagnetic systems, there are identical/non-identical opposite spin sublattices which we have denoted as A and B in this diagram. The compensated/uncompensated ferrimagnets can be considered as a deviation from the ideal antiferromagnets, while the compensated/uncompensated quasi-altermagnets can be considered as a deviation from the ideal altermagnets. Therefore, in the quasi-altermagnet picture, the spin up and spin down band dispersion can be random, there still exist blurred nodal planes and, on the either sides of these blurred nodal planes we observe the change in the sign of the spin splitting of the bands as in altermagnets. However, the strength of the spin splitting may not be equal on either side unlike the altermagnet. For regular ferrimagnets, there are no such restrictions. If we consider altermagnets as a subgroup of antiferromagnets, then the quasi-altermagnet can be considered as a subgroup of ferrimagnets.
Fig. 10 (a) shows the hexagonal MnTe supercell with a pair of non-magnetic atom substitutions, exhibiting SG $Pmm2$ (No. 25). The constant energy surfaces for the spin-up and spin-down channels, computed at E = EF – 0.7 eV, are shown in Fig. 10(c). These surfaces reveal a $g$-wave altermagnetic character, characterized by distinct yet symmetric features connected through a sixfold roto-invertion symmetry in each spin channel.
@@
Appendix B Effect of magnetic atom doping in MnTe
@@
A single substitution, as expected with different local spin moments at Fe site, breaks the rotational symmetry connecting the opposite spin sublattices, resulting in SG $P3m1$ (No. 156). The corresponding crystal structure and band structure are shown in Fig. 11 (a) and (d). The bands exhibit the characteristics of quasi-altermagnetism. In addition, new defect states are introduced, as shown in the band structure (Fig. 11 (d)).
@@
We further examined the pair substitution cases, where two Mn atoms with opposite spin moments are substituted by Fe. Similar to the pair of non-magnetic atom substitutions discussed in the main text, a pair of magnetic atom substitutions also leads to multiple configurations with different SG. When the dopant positions are arranged in such a way that mirror symmetry (Mz) of the system is preserved, then the resulting structure belongs to SG $P\bar{6}m2$ (No. 187) and will exhibit altermagnetic characteristics. The other configurations, where the proper or improper rotational symmetry is absent, lead to quasi-altermagnetism. Compared to non-magnetic atom substitution the probability of retaining altermagnetism is less when multiple magnetic atoms are substituted, since the magnetic neutrality is also affected. The other cases will result into quasi-altermagnetic conditions. Band structures displaying quasi-altermagnetic characteristics and ideal altermagnetic characteristics are shown in Fig. 11 (e) and (f). These results demonstrate that the quasi-altermagnetism discussed in the main text for non-magnetic atom doping is not restricted to such cases; it can also occur when magnetic atoms are introduced.
Appendix C Band diagram representation of different collinear magnetic structures
In this section, we have schematically illustrated the band diagram (see Fig. 12) of different classes of collinear magnetic materials, namely ferromagnetic metal (e.g., Fe3Sn2 [29]), ferromagnetic half-metal (e.g., CO2MnGa [12], ferromagnetic insulator (e.g., monolayer CrI3 [14], antiferromagnetic metal (e.g., FeSe  [31]), antiferromagnetic insulator (e.g., MnO [13]), uncompensated ferrimagnetic insulator (e.g., Y3Fe5O12 [28], compensated ferrimagnetic insulator (e.g., NiICl [30], compensated ferrimagnetic half-metal YI2 [30], compensated ferrimagnetic semi-metal (Mn3Al [21]), altermagnet (e.g., MnTe [10]), compensated quasi-altermagnet [e.g., Se doped MnTe (this work)], and uncompensated quasi-altermagnet [e.g., Sb and I doped MnTe (this work)].
Table 1: Symmetry operations and AHC in doped MnTe. SG, SSG [22], symmetry operations connecting same and opposite spin sublattices, MSG and the allowed AHC components for out-of-plane and in-plane Néel vector orientations (30° from $a$-axis) for different doped MnTe systems.
SG SSG Symmetry Operations Néel vector out-of-plane Néel vector in-plane
Connecting same spin sublattices Connecting opposite spin sublattices MSG AHC component MSG AHC component
Altermagnet
[…]
$Fmm2$ (42) $F^{-1}m^{-1}m^{1}2^{\infty m}1$ E, C2z Mx, My $Fm^{\prime}m^{\prime}2$ (42.222) $\sigma_{xy}$ $Fm^{\prime}m2^{\prime}$ (42.221) $\sigma_{xz}$
CmCm (8.32) $\sigma_{xz}$
$Pmm2$ (25) $P^{-1}m^{-1}m^{1}2^{\infty m}1$ E, C2z Mx, My $Pm^{\prime}m^{\prime}2$ (25.60) $\sigma_{xy}$ $Pm^{\prime}m2^{\prime}$ (25.59) $\sigma_{zx}$
PmPm (6.18) $\sigma_{xy}$
Quasi-altermagnet
[…]
$C2/m$ (12) $C^{1}2/^{1}m^{\infty m}1$ E, I, C2y, My none $C2^{\prime}/m^{\prime}$ (12.62) $\sigma_{xy},\sigma_{yz}$ $C2^{\prime}/m^{\prime}$ (12.62) $\sigma_{xy},\sigma_{yz}$
$P\bar{1}$ (2.4) $\sigma_{xy},\sigma_{yz},\sigma_{xz}$
Table 2: Numerical values of various hopping interaction strengths, antiferromagnetic Hund’s coupling, and onsite energies considered for tight-binding band structures (in eV unit). Here, $\Delta_{\text{Mn}}$ and $\Delta_{\text{Mn}^{\prime}}$ represent the strength of antiferromagnetic Hund’s coupling at the magnetic sites. $\alpha$ is the onsite energy of the magnetic atom, while $\beta_{1}$ and $\beta_{2}$ denote the effective onsite energies of the triangular nonmagnetic atom planes. $t_{M-T\sigma/\pi}$ ($t_{M-X\sigma/\pi}$) are $\sigma/\pi$ interaction strength between the Mn and Te (X=Sb/Se/I) atoms. $t_{T-T^{\prime}\sigma/\pi}$ ($t_{X-X^{\prime}\sigma/\pi}$) are $\sigma/\pi$ interactions between the nonequivalent nonmagnetic Te (X) atoms. $t_{T-X\sigma/\pi}$ are $\sigma/\pi$ interaction strength between the Te and X atoms.
Figs. Hund’s coupling Onsite energy Te-Te interaction Te-X interaction X-X interaction Mn-Te interaction Mn-X interaction
$\Delta_{\text{Mn}}$ $\Delta_{\text{Mn}^{\prime}}$ $\alpha$ $\beta_{1}$ $\beta_{2}$ $t_{T-T^{\prime}\sigma}$ $t_{T-T^{\prime}\pi}$ $t_{T-X\sigma}$ $t_{T-X\pi}$ $t_{X-X^{\prime}\sigma}$ $t_{X-X^{\prime}\pi}$ $t_{M-T\sigma}$ $t_{M-T\pi}$ $t_{M-X\sigma}$ $t_{M-X\pi}$
Altermagnet
8(d) 22 - 1.85-1.85 4.5-4.5 4.5-4.5 $1.112$ 0.325-0.325 $0.556$ 0.162-0.162 - - 1.42-1.42 $0.42$ 0.71-0.71 $0.21$
Quasi-altermagnet
9(d) 22 22 1.85-1.85 4-4 4.5-4.5 $1.112$ 0.325-0.325 $1.112$ 0.325-0.325 $1.112$ 0.325-0.325 1.42-1.42 $0.42$ 1.42-1.42 $0.42$
9(e) 22 22 1.85-1.85 4.5-4.5 4.5-4.5 $1.112$ 0.325-0.325 $0.889$ 0.26-0.26 $0.778$ 0.227-0.227 1.42-1.42 $0.42$ 1.14-1.14 $0.34$
9(f) $2.2$ 22 1.85-1.85 4.5-4.5 4.5-4.5 $1.112$ 0.325-0.325 $1.112$ 0.325-0.325 $1.112$ 0.325-0.325 1.42-1.42 $0.42$ 1.42-1.42 $0.42$
Appendix D Symmetry analysis and AHE in doped MnTe
Here we analyze the symmetries and the corresponding AHC components for all the doped cases. In the main text, the discussion primarily focused on systems with magnetization oriented along the out-of-plane direction. However, in the experimentally established MnTe structure, the Néel vector lies in the plane, and its orientation plays a crucial role in determining the AHC [17]. Therefore, this section is devoted to a detailed symmetry analysis- including both SSG and MSG-for different orientations of the Néel vector (in-plane and out-of-plane), for both undoped and doped systems.
Table 1 summarizes the SSG of all doped systems along with the symmetry operations connecting the same and opposite spin sublattices. From the analysis in the main text, we learned that in the single-dopant case the system consistently exhibits altermagnetic ordering. In contrast, the paired dopant scenario leads to several distinct configurations [see Fig. 3(e) of main text], with 46.67% maintaining altermagnetism, while the rest exhibit quasi-altermagnetism. As expected, in Table 1, we find that in altermagnets, opposite spin sublattices are related by rotational symmetry operations. In contrast quasi-altermagnets lack such symmetry operations leading to lack of compensation between the spin channels, thereby lifting the degeneracy between the up and down spin bands throughout the Brillouin zone.
The last two columns of Table 1 represent MSG and corresponding AHC components for both out-of-plane and in-plane orientation of the Néel vector. Detailed discussion for the out-of-plane Néel vector has been presented in section III.5. For all doped systems, an AHC is present when Néel vector lies in plane, similar to the pristine case [17]. Quasi-altermagnets exhibit multiple AHC components, whereas altermagnets exhibit single AHC component. For SG $Fmm2$ (No. 42), $Pmm2$ (No. 25), and $C2/m$ (No. 12), when the Néel vector lies in the plane, two distinct types of MSGs emerge, depending on the specific dopant position arrangements.
Appendix E Details of the tight binding Hamiltonian
The various interaction strengths, which are used to produce the tight-binding band structure, are shown in Figs. 8(d) and 9(d-f) of the main text are listed in Table 2. Note that, although a particular set of numerical values is used here to produce the band structure, the qualitative inferences made out of the tight-binding Hamiltonian remain invariant under the choice of the numerical values of the interaction strength and solely depend on the nature of the Hamiltonian of Eqs. 3 and 4 of the main text. For the case of altermagnetism in doped MnTe (case I), the Mn-Te and Mn-X interactions for the opposite spin sublattices are related via a mirror symmetry, leading to identical characteristic equations for up-spin and down-spin bands and thereby a perfect antiferromagnetic band structure. Te-Te and Te-X interactions lift this degeneracy by equal width and alter the spin characters of the bands in two consecutive antinodal zones, leading to altermagnetism. On the other hand, for the case of quasi-altermagnetism in doped MnTe (case II), there exists no symmetry relation between the hopping interactions of opposite spin sublattices. Therefore, unlike the former case, here the antiferromagnetic band degeneracy is lifted even in the absence of interactions among nonequivalent nonmagnetic atoms. Therefore, even if the alteration of spin characters happens in the consecutive antinodal zones, the splitting width is not identical, leading to quasi-altermagnetism.